OpenStudy (anonymous):

1976 squareoot (2a +b )1976

OpenStudy (anonymous):

1976 squareroot (2a +b ) ^ 1976 correction

OpenStudy (across):

\[\sqrt[1976]{2a+b}^{1976}\]Is this what you mean?

OpenStudy (anonymous):

yes but the 2a + b is in parentheses with the 1976 raised like this ^ 1976

OpenStudy (anonymous):

jus like dat

OpenStudy (across):

\[\sqrt[1976]{2a+b}^{1976}\]Can be rewritten as: \[(2a+b)^{\frac{1976}{1976}}\]The fraction simplifies to: \[(2a+b)^1\]Which is simply: \[2a+b\]

OpenStudy (anonymous):

thanks